OpenStudy (anonymous):

find the real soultions of the equation: 4(x+1)^2+14(x+1)+6=0

OpenStudy (anonymous):

Set n = x + 1 and rewrite the equation. Solve for "n" by factoring. Then go back to n = x + 1 to get "x".

OpenStudy (anonymous):

4(x+1)(x+1)+14(x+1)+6=0

OpenStudy (anonymous):

That way is too much work. Better to do it the way I described in my first post.

OpenStudy (anonymous):

not sure how

OpenStudy (anonymous):

4n^2 + 14n + 6 = 0 and solve for n. Much easier.

OpenStudy (anonymous):

i would subtract six from both sides

OpenStudy (anonymous):

No, that is also not the way to go. Just go the way I showed you. Solve for n. Once you get n, then n = x + 1. No subtracting 6 from each side. 4n^2 + 14n + 6 = 0 and solve for n.

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

Do you know how to solve quadratic equations by factoring?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Well, factoring is semi-guesswork anyway so don't worry too much. You can always solve by using the quadratic formula. But I'll factor the equation for you. 4n^2 + 14n + 6 = 0 -> (4n + 2)(n + 3) = 0 So, n = -1/2 or -3 And since n = x + 1 we have x = n - 1 So, x = -3/2 or -4 and that's your answer.